level curves and parametric equation
(1) Consider the function a, )1)( 2)2 (a) Find the level curves of /(x,y) for heights 0, 1 /2, 1, and 2, and plot them on the same 2D Aaph. Use that information, as well as any other information you think you midt need, tereketch the surface f(x,y).
(b) Find the parametric equation of the intersection of r2y4 with -f(r,y and sketch that parametric curve on the graph from part (a)
Answer #1

2 リー>の× ø ø ø (x-1)2 + (y-2)2=2 (x-1)2 + (y-2)2=1 (x-1)2 + (y-2)2= × × 2 0